OpenStudy (anonymous):

A tree casts a 24' long shadow. A bird sitting at the very top of the tree looking down at the shadow forms an angle of depression of 28°. Rounding to the nearest integer, how tall is the tree? ________ feet

OpenStudy (amistre64):

tan(28) = 24/h h = 24/tan(28) calculator says: 45.137

OpenStudy (amistre64):

thats off; I think....let me recheck that

OpenStudy (amistre64):

yeah, 28 is at the wrong end... tan(28) = h/24 24 tan(26) = h

OpenStudy (amistre64):

12.761 is correct

OpenStudy (amistre64):